Find the x-intercepts defined by the equation: y=3x^2+3x-6.
Answer by ewatrrr(24785)


Hi,
Finding the x-intercepts defined by the equation: y =3x^2+3x-6
3x^2 + 3x - 6 = 0 Note: y = 0 where the parabola crosses x-axis
3(x^2 + x - 2)= 0
Factoring
(x+2)(x-1)=0 Note:SUM of the inner product(2x) and the outer product(-x) = 1*x
(x+2)= 0
x = -2
(x-1)= 0
x = 1